OpenStudy (anonymous):

PLEASE HELP ME Find the first error that Tony made in his work and circle it. (there may be more than one mistake!) -6 ÷ 2/3 = -6 * 2/3 = -6/1 * 2/3 = 12/3 = 4

OpenStudy (anonymous):

would it be -6 ÷ 2/3 = -6 * 3/2 = -9?